# PeRFlow: Piecewise Rectified Flow as Universal Plug-and-Play Accelerator

### v1

@@ Abstract
We present Piecewise Rectified Flow (PeRFlow), a flow-based method for accelerating diffusion models. PeRFlow divides the sampling process of generative flows into several time windows and straightens the trajectories in each interval via the reflow operation, thereby approaching piecewise linear flows. PeRFlow achieves superior performance in a few-step generation. Moreover, through dedicated parameterizations, the obtained PeRFlow models show advantageous transfer ability, serving as universal plug-and-play accelerators that are compatible with various workflows based on the pre-trained diffusion models. The implementations of training and inference are fully open-sourced. [^1] https://github.com/magic-research/piecewise-rectified-flow
@@ 1 Introduction
Diffusion models can be efficiently sampled by solving the corresponding probability flow ordinary differential equations (PF-ODEs) [Song et al., 2021, 2022]. Researchers have designed many advanced samplers, such as DDIM [Song et al., 2022], DPM-solver [Lu et al., 2022b], and DEIS [Zhang and Chen, 2022], to accelerate generation, inspired by the semi-linear structure and adaptive solvers in ODEs. However, these samplers still require tens of inference steps to generate satisfying results. Researchers have also explored distilling pretrained diffusion models into few-step generative models [Salimans and Ho, 2022, Meng et al., 2023, Gu et al., 2023, Yin et al., 2023, Nguyen and Tran, 2024], which have succeeded in synthesizing images within 8 inference steps. Progressive Distillation [Salimans and Ho, 2022] separates the whole sampling process into multiple segments and learns the mapping from starting points to endpoints for each segment. Distribution Matching Distillation [Yin et al., 2023] and SwiftBrush [Nguyen and Tran, 2024] use the score distillation loss to align the distributions of teacher and one-step student generators. UFOGen [Xu et al., 2023a], SDXL-Turbo [Sauer et al., 2023] and SDXL-Lightning [Lin et al., 2024] resort to adversarial training for learning few-step/one-step image generators. They initialize the students from pretrained diffusion models and use adversarial and/or MSE losses to align the student model’s generation with the pretrained ones. These methods suffer from the difficult tuning of the adversarial training procedure and the mode collapse issue. Latent Consistency Model (LCM)  [Luo et al., 2023a, b] adopts consistency distillation [Song et al., 2023] to train a generator that directly maps noises to the terminal images. LCM only utilizes supervised distillation where the training procedure will be more stable and easier in comparison to adversarial training. However, the generated images have fewer details compared with SDXL-Lighting.
@@
Unlike the existing methods above, which mainly learn the mappings from noises to images, we aim to simplify the flow trajectories and preserve the continuous flow trajectories of the original pretrained diffusion models. Specifically, we attempt to straighten the trajectories of the original PF-ODEs via a piecewise reflow operation. Previously, InstaFlow [Liu et al., 2023] leverages the rectified flow framework[Liu et al., 2022, Liu, 2022] to learn the transformation from initial random noise to images. It bridges the two distributions with linear interpolation and trains the model by matching the interpolation. With the reflow operation, it may be able to learn straight-line flows for one-step generation via pure supervised learning. InstaFlow provides a simple pipeline for accelerating pretrained diffusion models, however, it suffers from poor sampling quality which can be attributed to synthetic data generation. The reflow operation requires generating data from the pretrained diffusion models with ODE solvers (e.g., DDIM or DPM-Solver [Lu et al., 2022b, 2023]) to construct a training dataset. Synthesizing training data brings two problems: (1) constructing and storing the dataset requires excessive time and space, which limits its training efficiency; (2) synthetic data has a noticeable gap with real training data in quality due to the numerical error of solving ODEs. Thus, the performance of the learned straighter flow is bounded. To address the problems, we propose piecewise rectified flow (PeRFlow), which divides the flow trajectories into several time windows and conducts reflow in each window. By solving the ODEs in the shortened time interval, PeRFlow avoids simulating the entire ODE trajectory for preparing the training data. This significantly reduces the target synthesis time, enabling the simulation to be performed in real time along with the training procedure. Besides, PeRFlow samples the starting noises by adding random noises to clean images according to the marginal distributions, and solves the endpoints of a shorter time interval, which has a lower numerical error than integrating the entire trajectories. Through such a divide-and-conquer strategy, PeRFlow can straighten the sampling trajectories with large-scale real training data. Besides the training framework, we also design a dedicated parameterization method for PeRFlow to inherit sufficient knowledge from the pretrained diffusion models. Diffusion models are usually trained with $\epsilon$-prediction, but flow-based generative models generate data by following the velocity field. We derive the correspondence between $\epsilon$-prediction and the velocity field of flow, thus narrowing the gap between the pretrained diffusion models and the student PeRFlow model. Consequently, PeRFlow acceleration converges fast and the resultant model can synthesize highly-detailed images within very few steps. PeRFlow does not require unstable adversarial training or a complete modification of the training paradigm. It is a lightweight acceleration framework and can be easily applied to training unconditional/conditional generative models of different data modalities.
@@ 2.1 Rectified Flow and Reflow
A favorable property of the reflow operation is that, by using the optimal $v_{\theta^{*}}$ learned in (2) as teacher, one can repeat the reflow operation to learn a new flow-based generative model which is straighter and faster in simulation,
$\min_{\theta^{\prime}}\mathbb{E}_{{\bm{z}}_{1}\sim\pi_{1}}$ $\left[\int_{0}^{1}\|({\bm{z}}_{1}-{\bm{z}}_{0})-v_{\theta^{\prime}}({\bm{z}}_{t},t)\|^{2}\mathrm{d}t\right],$ (3)
$\text{with}\quad{\bm{z}}_{0}={\bm{z}}_{1}+\int_{1}^{0}$ $v_{\theta^{*}}({\bm{z}}_{t},t)\mathrm{d}t\quad\text{and}\quad{\bm{z}}_{t}=(1-t){\bm{z}}_{0}+t{\bm{z}}_{1}.$
Pretrained diffusion models can be transformed to probability flow ODEs, hence they can serve as $v_{\theta^{*}}$ for learning the straighter flow in (3). InstaFlow [Liu et al., 2023] proposed to accelerate pretrained diffusion-based text-to-image models via reflow. Given a pretrained diffusion model $f_{\phi}$, one can generate new data by solving the corresponding probability flow ODE. We denote $\Phi({\bm{z}}_{t},t,s)$ as the ODE solver, such as the DPM-Solver [Lu et al., 2022a]. For simplicity, our notation drops the parameters in the ODE solvers. By simulating with $\Phi({\bm{z}}_{1},1,0)$, where ${\bm{z}}_{1}$ is sampled from the random Gaussian distribution $\pi_{1}$, it synthesizes a dataset of (text, noise, image) triplets for reflow. Since it usually takes tens of inference steps to generate high-quality data with $\Phi({\bm{z}}_{1},1,0)$, InstaFlow is expensive to scale up. Moreover, since InstaFlow is trained with generated images, it lacks the supervision of real data and thus compromises the resulting generation quality. In the following subsections, we target solving these problems.
@@ 2.2 Piecewise Rectified Flow
Scaling Up with Real Training Data
PeRFlow divides the time range $[1,0]$ into multiple windows. For each window, the starting point ${\bm{z}}_{t_{k}}$ is obtained by adding random noise to real training data $z_{0}$, and it only requires several inference steps to solve the ending point ${\bm{z}}_{t_{k-1}}$. The computational cost is significantly reduced for each training iteration compared to InstaFlow, allowing us to train PeRFlow on large-scale training datasets with fast online simulation of the ODE trajectory. Besides, solving endpoints of a shorter time window $[{\bm{z}}_{t_{k}},{\bm{z}}_{t_{k-1}})$ has lower numerical errors in comparison to the entire time range. High-quality supervision yields significant improvement in the generation results.
Classifier-Free Guidance in Training
Classifier-free guidance (CFG) [Ho and Salimans, 2021] is a common technique to improve the generation quality of text-to-image models. During training, we solve the endpoints ${\bm{z}}_{t_{k-1}}$ for each time window $[{t_{k}},{t_{k-1}})$ in an online manner via an ODE solver $\Phi({\bm{z}}_{t_{k}},t_{k},t_{k-1},{\bm{c}},w)$, where $w\geq 1$ denotes the CFG scale, ${\bm{c}}$ denotes the text prompt. CFG is turned off when $w=1$. PeRFlow supports two modes: CFG-sync and CFG-fixed:
PeRFlow as Universal Plug-and-Play Accelerator
PeRFlow initializes the weights of the student model $\theta$ with the pretrained diffusion model $\phi$. After training with piecewise reflow, we find that the change of weights $\Delta W=\theta-\phi$ can be used to seamlessly accelerate many other workflows pretrained with the diffusion model. For exmaple, $\Delta W$ of PeRFlow+SD-v1.5 can accelerate the ControlNets [Zhang and Agrawala, 2023], IP-Adaptor [Ye et al., 2023] and multiview generation [Long et al., 2023] pipelines trained with the original SD v1.5. The accelerated pipelines achieve nearly lossless few-step generation as the original many-step generation. Please refer to Section 3.3 for detailed results.
@@ 9      Solve the endpoint of the time window $z_{t_{k-1}}=\Phi({\bm{z}}_{t_{k}},t_{k},t_{k-1})$ ;
21                Return: Fast PeRFlow $f_{\theta}$ and $\Delta W$.
@@ 3 Experiments
Experiment Configuration
All the PeRFlow models are initialized from their diffusion teachers. PeRFlow-SD-v1.5 is trained with images in resolution of $512\times 512$ using $\epsilon$-prediction defined in (8). PeRFlow-SD-v2.1 is trained with images in resolution of $768\times 768$ using ${\bm{v}}$-prediction. PeRFlow-SDXL is trained with images in resolution of $1024\times 1024$ using $\epsilon$-prediction defined in (9). Images are all sampled from the LAION-Aesthetics-5+ dataset [Schuhmann et al., 2022] and center-cropped. We also train PeRFlow-AnimateDiff with video clips in size of $16\times 384\times 384$ using $\epsilon$-prediction defined in (9). We randomly drop out the text captions with a low probability ($10\%$) to enable classifier-free guidance during sampling. We divide the time range $[0,1]$ into four windows uniformly. For each window, we use the DDIM solver to solve the endpoints with 8 steps. We refer to the Hugging Face scripts for training Stable Diffusion  [^2] https://github.com/huggingface/diffusers/tree/main/examples/text_to_image to set other hyper-parameters, including learning rate and weight decay.
@@ 3.1 Few-step generation
PeRFlow succeeds in accelerating pretrained Stable Diffusion models to few-step generators. As shown in figure 4, 5, and 3, PeRFlow can generate astonishing pictures with only 4 steps. If increasing the number of inference steps (e.g., 6 and 8), we can obtain images with much richer details as shown in figure 7. We compare the generation results with two SOTA methods, LCM-LORA and SDXL-lightning. PeRFlow enjoys richer visual details, refer to figure 2.
We compute the FID values of PeRFlow-SD-v1.5 in table 1. We randomly sample 30k images from LAION-5B-Aesthetics as the reference set and generate the same number of images via the PeRFlow model with 4/8 steps as the test sets. We also compute the FID values by using COCO2014 as the reference distributions. In comparison to LCM-LORA, we observe that PeRFlow models have obviously lower FID values. When increasing the number of inference steps, the FID values of PeRFlow will decrease because the numerical errors of solving sampling ODE are better controlled. However, the FID values of LCM-LORA unexpectedly increase.
@@
FID 4-step 8-step 4-step 8-step 4-step 8-step
PeRFlow 9.74 8.62 11.31 14.16 9.46 5.05
@@
Table 1: FID values of PeRFlow-SD-v1.5
(a) LCM-LORA for SDXL
(b) SDXL-lightning
(c) PeRFlow for SDXL
Figure 2: Images ($1024\times 1024$) generated by PeRFlow enjoy richer details and texture in comparison to other acceleration methods. Prompt-1: “A closeup face photo of girl, wearing a raincoat, in the street, heavy rain, bokeh”; Prompt-2: “A closeup face photo of a boy in white shirt standing on the grassland, flowers”; Prompt-3: “A beautiful cat bask in the sun”.
Figure 3: 4-step sampling ($1024\times 1024$) via PeRFlow-SDXL.
Figure 4: 4-step sampling ($512\times 512$) via PeRFlow-SD-v1.5.
Figure 5: 4-step sampling ($768\times 768$) via PeRFlow-SD-v2.1.
Figure 6: 6-step sampling ($16\times 512\times 512$) via PeRFlow-AnimateDiff (motion module-v3 with DreamShaper as the 2D base). The text prompts used are “A young woman smiling, in the park, sunshine” and “A dog sitting in the garden, snow, trees”. We append “4k uhd, high quality, masterpiece” as the prefix of prompts, and use “distorted, blur, haze, warm, over-saturated, worst quality, low quality, letterboxed” as the negative prompt.
Figure 7: Comparison between 4-step (left) and 5-step (right) generation of PeRFlow-SDXL.
3.2 Domain shift caused by acceleration
Diffusion acceleration aims to speed up the sampling process of the pretrained models and simultaneously preserves their performance and properties. We compute the FID values between the generation of the original SD-v1.5 and the acceleration methods. From table 1, we observe the FID values of PeRFlow are smaller than LCM-LORA. If using more inference steps, the FID values of PeRFlow will further drop. This implies the distribution shift to the original SD caused by PeRFlow is much smaller than LCM-LORA. The numerical comparison corresponds to the results in figure 8. The color style and layout of PeRFlow’s results match the results of the pretrained models, while an obvious domain shift appears in the results of LCM-LORA. Besides, the sampling diversity of PeRFlow is similar to the original SD-v1.5 and looks better than LCM-LORA in figure 9.
@@ 3.2 Domain shift caused by acceleration
Figure 9: PeRFlow has better sampling diversity compared to LCM-LORA.
3.3 Universal acceleration to SD workflows

### v3

@@ Abstract
We present Piecewise Rectified Flow (PeRFlow), a flow-based method for accelerating diffusion models. PeRFlow divides the sampling process of generative flows into several time windows and straightens the trajectories in each interval via the reflow operation, thereby approaching piecewise linear flows. PeRFlow achieves superior performance in a few-step generation. Moreover, through dedicated parameterizations, the PeRFlow models inherit knowledge from the pretrained diffusion models. Thus, the training converges fast and the obtained models show advantageous transfer ability, serving as universal plug-and-play accelerators that are compatible with various workflows based on the pre-trained diffusion models. Codes for training and inference are publicly released.  [^1] https://github.com/magic-research/piecewise-rectified-flow.
@@ 1 Introduction
Diffusion models can be efficiently sampled by solving the corresponding probability flow ordinary differential equations (PF-ODEs) [Song et al., 2021, 2022]. Researchers have designed many advanced samplers, such as DDIM [Song et al., 2022], DPM-solver [Lu et al., 2022b], and DEIS [Zhang and Chen, 2022], to accelerate generation, inspired by the semi-linear structure and adaptive solvers in ODEs. However, these samplers still require tens of inference steps to generate satisfying results. Researchers have also explored distilling pretrained diffusion models into few-step generative models [Salimans and Ho, 2022, Meng et al., 2023, Gu et al., 2023, Yin et al., 2023, Nguyen and Tran, 2024, Berthelot et al., 2023], which have succeeded in synthesizing images within 8 inference steps. Progressive Distillation [Salimans and Ho, 2022] separates the whole sampling process into multiple segments and learns the mapping from starting points to endpoints for each segment. Distribution Matching Distillation [Yin et al., 2023] and SwiftBrush [Nguyen and Tran, 2024] use the score distillation loss to align the distributions of teacher and one-step student generators. UFOGen [Xu et al., 2023a], SDXL-Turbo [Sauer et al., 2023] and SDXL-Lightning [Lin et al., 2024] resort to adversarial training for learning few-step/one-step image generators. They initialize the students from pretrained diffusion models and use adversarial and/or MSE losses to align the student model’s generation with the pretrained ones. These methods suffer from the difficult tuning of the adversarial training procedure and the mode collapse issue. Latent Consistency Model (LCM)  [Luo et al., 2023a, b] adopts consistency distillation [Song et al., 2023] to train a generator that directly maps noises to the terminal images. LCM only utilizes supervised distillation where the training procedure will be more stable and easier in comparison to adversarial training. However, the generated images have fewer details compared with SDXL-Lighting.
@@
Unlike the existing methods above, which mainly learn the mappings from noises to images, we aim to simplify the flow trajectories and preserve the continuous flow trajectories of the original pretrained diffusion models. Specifically, we attempt to straighten the trajectories of the original PF-ODEs via a piecewise reflow operation. Previously, InstaFlow [Liu et al., 2023] leverages the rectified flow framework[Liu et al., 2022, Liu, 2022] to learn the transformation from initial random noise to images. It bridges the two distributions with linear interpolation and trains the model by matching the interpolation. With the reflow operation, it may be able to learn straight-line flows for one-step generation via pure supervised learning. InstaFlow provides a simple pipeline for accelerating pretrained diffusion models, however, it suffers from poor sampling quality which can be attributed to synthetic data generation. The reflow operation requires generating data from the pretrained diffusion models with ODE solvers (e.g., DDIM or DPM-Solver [Lu et al., 2022b, 2023]) to construct a training dataset. Synthesizing training data brings two problems: (1) constructing and storing the dataset requires excessive time and space, which limits its training efficiency; (2) synthetic data has a noticeable gap with real training data in quality due to the numerical error of solving ODEs. Thus, the performance of the learned straighter flow is bounded.
To address the problems, we propose piecewise rectified flow (PeRFlow), which divides the flow trajectories into several time windows and conducts reflow in each window. By solving the ODEs in the shortened time interval, PeRFlow avoids simulating the entire ODE trajectory for preparing the training data. This significantly reduces the target synthesis time, enabling the simulation to be performed in real time along with the training procedure. Besides, PeRFlow samples the starting noises by adding random noises to clean images according to the marginal distributions, and solves the endpoints of a shorter time interval, which has a lower numerical error than integrating the entire trajectories. Through such a divide-and-conquer strategy, PeRFlow can straighten the sampling trajectories with large-scale real training data. Besides the training framework, we also design a dedicated parameterization method for PeRFlow to inherit sufficient knowledge from the pretrained diffusion models. Diffusion models are usually trained with $\epsilon$-prediction, but flow-based generative models generate data by following the velocity field. We derive the correspondence between $\epsilon$-prediction and the velocity field of flow, thus narrowing the gap between the pretrained diffusion models and the student PeRFlow model. Consequently, PeRFlow acceleration converges fast and the resultant model can synthesize highly-detailed images within very few steps. PeRFlow does not require unstable adversarial training or a complete modification of the training paradigm. It is a lightweight acceleration framework and can be easily applied to training unconditional/conditional generative models of different data modalities.
@@ 2.1 Rectified Flow and Reflow
InstaFlow [Liu et al., 2023] proposed to accelerate pretrained diffusion-based text-to-image models via reflow. Given a pretrained diffusion model $f_{\phi}$, one can generate new data by solving the corresponding probability flow ODE. We denote $\Phi({\bm{z}}_{t},t,s)$ as the ODE solver, such as the DPM-Solver [Lu et al., 2022a]. For simplicity, our notation drops the parameters in the ODE solvers. By simulating with ${\bm{z}}_{0}=\Phi({\bm{z}}_{1},1,0)$, where ${\bm{z}}_{1}$ is sampled from the random Gaussian distribution $\pi_{1}$, it synthesizes a dataset of (text, noise, image) triplets for reflow. Since it usually takes tens of inference steps to generate high-quality data with $\Phi({\bm{z}}_{1},1,0)$, InstaFlow is expensive to scale up. Moreover, since InstaFlow is trained with generated images, it lacks the supervision of real data and thus compromises the resulting generation quality. In the following subsections, we target solving these problems.
@@ 9      Solve the endpoint of the time window $z_{t_{k-1}}=\Phi({\bm{z}}_{t_{k}},t_{k},t_{k-1})$ ;
Return: Fast PeRFlow $f_{\theta}$ and $\Delta W$.
@@
Scaling Up with Real Training Data   PeRFlow divides the time range $[1,0]$ into multiple windows. For each window, the starting point ${\bm{z}}_{t_{k}}$ is obtained by adding random noise to real training data $z_{0}$, and it only requires several inference steps to solve the ending point ${\bm{z}}_{t_{k-1}}$. The computational cost is significantly reduced for each training iteration compared to InstaFlow, allowing us to train PeRFlow on large-scale training datasets with fast online simulation of the ODE trajectory. Besides, solving endpoints of a shorter time window $[{\bm{z}}_{t_{k}},{\bm{z}}_{t_{k-1}})$ has lower numerical errors in comparison to the entire time range. High-quality supervision yields significant improvement in the generation results.
Classifier-Free Guidance in Training   Classifier-free guidance (CFG) [Ho and Salimans, 2021] is a common technique to improve the generation quality of text-to-image models. During training, we solve the endpoints ${\bm{z}}_{t_{k-1}}$ for each time window $[{t_{k}},{t_{k-1}})$ in an online manner via an ODE solver $\Phi({\bm{z}}_{t_{k}},t_{k},t_{k-1},{\bm{c}},w)$, where $w\geq 1$ denotes the CFG scale, ${\bm{c}}$ denotes the text prompt. CFG is turned off when $w=1$. PeRFlow supports two modes: CFG-sync and CFG-fixed:
PeRFlow as Universal Plug-and-Play Accelerator   PeRFlow initializes the weights of the student model $\theta$ with the pretrained diffusion model $\phi$. After training with piecewise reflow, we find that the change of weights $\Delta W=\theta-\phi$ can be used to seamlessly accelerate many other workflows pretrained with the diffusion model. For exmaple, $\Delta W$ of PeRFlow+SD-v1.5 can accelerate the ControlNets [Zhang and Agrawala, 2023], IP-Adaptor [Ye et al., 2023] and multiview generation [Long et al., 2023] pipelines trained with the original SD v1.5. The accelerated pipelines achieve nearly lossless few-step generation as the original many-step generation. Please refer to Section 3.2 for detailed results.
@@ 3 Experiments
Experiment Configuration   All the PeRFlow models are initialized from their diffusion teachers. PeRFlow-SD-v1.5 is trained with images in resolution of $512\times 512$ using $\epsilon$-prediction defined in (7). PeRFlow-SD-v2.1 is trained with images in resolution of $768\times 768$ using ${\bm{v}}$-prediction. PeRFlow-SDXL is trained with images in resolution of $1024\times 1024$ using $\epsilon$-prediction defined in (8). Images are all sampled from the LAION-Aesthetics-5+ dataset [Schuhmann et al., 2022] and center-cropped. We also train PeRFlow-AnimateDiff with video clips in size of $16\times 384\times 384$ using $\epsilon$-prediction defined in (8). We randomly drop out the text captions with a low probability ($10\%$) to enable classifier-free guidance during sampling. We divide the time range $[0,1]$ into four windows uniformly. For each window, we use the DDIM solver to solve the endpoints with 8 steps. We refer to the Hugging Face scripts for training Stable Diffusion  [^2] https://github.com/huggingface/diffusers/tree/main/examples/text_to_image to set other hyper-parameters, including learning rate and weight decay. All experiments are conducted with $16$ NVIDIA A100 GPUs.
@@ 3.1 Few-step generation
PeRFlow succeeds in accelerating pretrained Stable Diffusion models to few-step generators. As shown in figure 2 and 3, PeRFlow can generate astonishing pictures with only 4 steps. If increasing the number of inference steps (e.g., 5 or 6), we can obtain images with much richer details. We compare the generation results with recent acceleration methods, including InstaFlow, LCM-LORA, and SDXL-lightning. PeRFlow enjoys richer visual texture and better alignment between text prompts and images. Refer to figure 10, 12, and 13 in Appendix for more results.
We compute the FID values of PeRFlow-accelerated SDs in table 1 using images on three different reference distributions: (1) LAION-5B-Aesthetics [Schuhmann et al., 2022], which is the training set of PeRFlow and other methods; (2) MS COCO 2014 [Lin et al., 2014] validation dataset; (3) images generated from SD- v1.5/XL with JourneyDB [Sun et al., 2024] prompts. We generate 30,000 images for the SD-v1.5 models and 10,000 for the SDXL series. We set the inference steps to 4 and 8 steps, respectively. In comparison to LCM-LoRA, we observe that PeRFlow models have obviously lower FID values. When increasing the number of inference steps, FID values of PeRFlow decrease because the numerical errors of solving ODE are better controlled. However, FID values of LCM-LoRA unexpectedly increase.
Figure 2: The $1024\times 1024$ images generated by PeRFlow enjoy richer details and better text-image consistency in comparison to other acceleration methods on SDXL. Prompt #1: “a closeup face photo of girl, wearing a raincoat, in the street, heavy rain, bokeh”; Prompt #2: “a closeup face photo of a boy in white shirt standing on the grassland, flowers”; Prompt #3: “a huge red apple in front of a small dog, heavy snow”. Prompt #4: “front view of a boat sailing in a cup of water”.
Figure 3: The $512\times 512$ images generated by PeRFlow enjoy richer details and color styles in comparison to other acceleration methods on SD-v1.5 (w/ DreamShaper). Images in each row are generated with the same random seed.
Figure 4: 6-step generation ($16\times 512\times 512$) via PeRFlow-AnimateDiff (motion module-v3 with DreamShaper). The text prompts used are “A young woman smiling, in the park, sunshine” and “A dog sitting in the garden, snow, trees”.
@@
Method 4-step 8-step 4-step 8-step 4-step 8-step
InstaFlow 14.32 10.98 13.86 11.40 16.67 10.45
@@
PeRFlow 8.60 8.52 11.31 14.16 8.28 5.03
(a) SD-v1.5
LAION-5B COCO2014 SDXL
Method 4-step 8-step 4-step 8-step 4-step 8-step
Lightning 15.47 14.37 22.86 20.44 11.41 10.49
LCM-LoRA 13.66 13.31 19.74 21.70 9.42 9.90
PeRFlow 13.30 13.06 18.48 19.21 9.28 9.12
(b) SDXL
Table 1: FID values of different acceleration methods.
Domain shift caused by acceleration   When accelerating diffusion models, we expect to preserve the performance and properties of the pretrained models. In table 1, we compute the FID values between the generation of the original SD models and the accelerated models. We observe the FID values of PeRFlow are smaller than LCM-LORA, InstaFlow, and SDXL-Lightning. This implies the distribution shift to the original SD models caused by PeRFlow is much smaller than other counterparts. The numerical comparison corresponds to the results in figure 5. The color style and layout of PeRFlow’s results match the results of the pretrained models, while an obvious domain shift appears in the results of LCM-LoRA. Besides, the sampling diversity of PeRFlow is similar to the original SD-v1.5 and appears to be better than LCM-LoRA in figure 6.
@@
Figure 6: Three random samples from two models with the same prompts. PeRFlow has better sampling diversity compared to LCM-LoRA.
3.2 PeRFlow as Universal Plug-and-Play Accelerator on SD Work Flows